# TrajAgent: An Agent Framework for Unified Trajectory Modelling

### v1

@@ Abstract.
Trajectory modeling, which includes research on trajectory data pattern mining and future prediction, has widespread applications in areas such as life services, urban transportation, and public administration. Numerous methods have been proposed to address specific problems within trajectory modelling. However, due to the heterogeneity of data and the diversity of trajectory tasks, achieving unified trajectory modelling remains an important yet challenging task. In this paper, we propose TrajAgent, a large language model-based agentic framework, to unify various trajectory modelling tasks. In TrajAgent, we first develop UniEnv, an execution environment with a unified data and model interface, to support the execution and training of various models. Building on UniEnv, we introduce TAgent, an agentic workflow designed for automatic trajectory modelling across various trajectory tasks. Specifically, we design AutOpt, a systematic optimization module within TAgent, to further improve the performance of the integrated model. With diverse trajectory tasks input in natural language, TrajAgent automatically generates competitive results via training and executing appropriate models. Extensive experiments on four tasks using four real-world datasets demonstrate the effectiveness of TrajAgent in unified trajectory modelling, achieving an average performance improvement of 15.43% over baseline methods.
@@
copyright: acmcopyrightcopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
With the rapid development of web services and mobile devices (Zheng, 2015; Chen et al., 2024), large-scale trajectory data, such as check-in data from social network (Yang et al., 2016), have been collected, greatly facilitating research in trajectory modelling. Trajectory modelling involves the processing, mining and prediction of trajectory data, with widespread applications in urban transportation, location services and public management. The typical areas of trajectory modelling (Chen et al., 2024; Graser et al., 2024) can be classified into three main categories: trajectory classification (Liang et al., 2022), trajectory prediction (Yang et al., 2022), and trajectory generation (Wei et al., 2024). Each category encompasses various sub-tasks; for instance, the trajectory prediction task can be further divided into next location prediction task (Liu et al., 2016), final destination prediction task (Zhang et al., 2019), and travel time estimation task (Wang et al., 2018), among others. Given the huge value of trajectory modelling in diverse practical applications, various algorithms and models (Graser et al., 2024) have been proposed to address these tasks, particularly deep learning-based models in recent years. This has facilitated significant advancements in the field, with many tasks achieving a high level of performance.
@@
However, existing methods are designed for specific tasks and datasets, making it difficult to share them across different tasks and data sources. For example, TrajFormer (Liang et al., 2022) is tailored for trajectory classification and cannot be applied in trajectory prediction or trajectory generation. Flash-back (Yang et al., 2020) is designed for sparse check-in trajectory prediction and is not suitable for dense GPS trajectory or road network-based trajectory modelling. In other words, due to the heterogeneity of application scenario and the diverse nature of trajectory data–varying in resolution, format and geographical regions–existing methods can only be applied in limited task with specific data for specific regions. Without a unified model, the deployment and application of these advanced models become costly and inefficient, hindering further development of trajectory modelling area. Therefore, developing a unified model for various trajectory modelling tasks is both emergent and fraught with significant challenges.
In last two years, the rapid development of large language models (LLMs) (OpenAI, 2024; Touvron et al., 2023) with extensive commonsense and powerful reasoning abilities presents great potential for solving challenging tasks, such as mathematics (Romera-Paredes et al., 2024) and logic games (Yao et al., 2022b). Furthermore, LLM based agent (Xi et al., 2023; Sumers et al., 2023) have emerged as a new paradigm, offering more robust and intelligent task-solving capabilities. These agents have demonstrated stable and promising performance in tackling complex real-world tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024). Leveraging the powerful understanding and reasoning abilities of LLM based agents, researchers explore the potential of utilizing LLM based agent for the unified modelling of diverse machine learning tasks (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024). For example, HuggingGPT (Shen et al., 2024) utilize LLM as a core manager to address various machine learning tasks with existing AI models, VisionLLM (Wu et al., 2024) investigates unified modelling for vision tasks across different domains, while Mora (Yuan et al., 2024) designs a multi-agent framework for replicating Sora’s functionalities with combining existing open source models.
Inspired by these works, we explore the potential of addressing the unified trajectory modeling problem using an LLM-based agent framework. However, designing an LLM-based agent for unified trajectory modelling is challenging due to several factors. Firstly, trajectory data and existing models for various trajectory modelling tasks are diverse and often inconsistent, making the integration and deployment of these models difficult. Secondly, the processing steps from raw trajectory data to final output results are lengthy and cumbersome (Chen et al., 2024), resulting in a large action space for the planning and execution of entire process. Third, the performance of existing models highly depends one data adaptation and parameter optimization, making the direct usage of existing pre-trained models infeasible.
Figure 1. Framework of TrajAgent, including UniEnv with unified data and model interface for running experiment, TAgent an agentic workflow for executing the whole trajectory modeling procedure, and AutOpt for the model enhancement and adaptation on selected data.
@@
In this paper, we propose TrajAgent, a systematic agent framework for unified trajectory modelling and analytics. Within TrajAgent, we first design a unified environment, UniEnv, for processing diverse trajectory data and preparing a cohesive running environment for various trajectory modelling tasks. In UniEnv, we define unified data and model interfaces to support the easy execution of various methods for trajectory modelling. Next, we develop an agentic workflow called TAgent for the automatic planning and execution of trajectory modelling tasks. In TAgent, we decompose the task workflow into four steps: task understanding, task planning, task execution and task summarization. In each step, we design a specialized expert agent to perform specific operations. Notably, we introduce AutOpt to enhance the task execution module, enabling the adaption of model performance through various automatic optimization mechanisms, including data augmentation, parameter optimization and joint optimization. In summary, our contributions are as follows,
@@
  • To the best of our knowledge, TrajAgent is the first attempt to employ an LLM-based agentic framework for unified trajectory modeling across diverse trajectory data and various trajectory tasks.
  • In TrajAgent, we design a agentic workflow to decompose the whole trajectory modelling task into several sub-tasks, all carried out within a unified environment UniEnv. Finally, we develop a comprehensive model optimization module to further enhance the model performance on targeted data.
@@
  • Extensive experiments on four representative trajectory modelling tasks across four real-world datasets demonstrate the effectiveness of proposed framework, which the optimized model achieving an average performance gain of 15.43%.
@@ 2. Methods
Figure 1 presents the whole framework of TrajAgent. It contains three components: 1) UniEnv, an environment with a unified data and model interface that supports the execution and training of various models; 2) TAgent, an agentic workflow designed to automatically complete diverse trajectory modelling task. It accepts requirements in the form of natural language input and directly outputs the adapted models; 3) AutOpt, an additional optimization module for TAgent that enhances the performance of specific model through data and parameter adaptation. Details of each component are presented as follows.
2.1. UniEnv
@@ 2.1.1. Task Interface
We summarize the supported trajectory modelling tasks in UniEnv as follows.
  • Trajectory prediction task: next location prediction and travel time estimation.
  • Trajectory classification task: trajectory user linking and trajectory anomaly detection.
  • Trajectory generation task: trajectory recovery and trajectory generation.
To enhance the clarity and effectiveness of understanding users’ language queries, we provide a detailed language description for each task. This description helps extract precise task requirements from user queries and facilitates subsequent data and model selection processes.
@@ 2.1.3. Model Interface
To support various model training in TrajAgent, we select at least one well-known model for each task and adapt them to match the running environments in UniEnv. Furthermore, we extract the semantic context from original paper of each model with the txyz.ai APIs  [^1] https://www.txyz.ai/ to generate detailed description of model. In this description, the verified data information and supported trajectory task information are recorded which supports the data and model selection in the subsequent trajectory planning stage of TrajAgent.
@@ 2.1.4. External Tools
To extend the capabilities of TrajAgent, we collect several external tools in UniEnv, including paper context extraction tool txyz.ai, hyperparameter optimization tool optuna  [^2] https://optuna.org/, processing and visualization tool for trajectory data movingpandas  [^3] https://github.com/movingpandas/movingpandas, open street map data processing tool osmnx  [^4] https://github.com/gboeing/osmnx.
Here, we also regard the LLM APIs utilized in the agentic workflow as one of the interface in the UniEnv, including the ChatGPT API  [^5] https://openai.com/ and DeepInfra  [^6] https://deepinfra.com for open-source LLMs.
2.2. TAgent
As shown in the middle of Figure 1, the architecture of TAgent is organized into four key modules: Task Understanding, Task Planning, Task Optimization, and Task Summary, which form an automated processing chain from user query to final result, eliminating the need for human-in-the-loop. Specifically, the task understanding module first receives user instructions in natural language form, and analyzes and identifies the type, name, and other key information of the tasks involved. Then, the task planning module will plan for the identified task, including dataset matching and model selection. Next, the task execution module executes the planning task and cooperate with the additional model optimization module AutOpt to further improve the task performance from both data augmentation and parameter selection perspectives. Last, the task summary module generate an analytical report of the task based on historical interactions and decisions of TAgent. Following the common practice(Shinn et al., 2024), each module in TAgent can be regarded as a small agent, which consists of a function module for executing its core function, a memory for recording the history interaction, and a reflection module for learning practice from the memory.
@@ 2.2.1. Task Understanding Module
As the first module of TAgent, task understanding module is designed to interact with user and extract detailed task information to launch subsequent stages. Given the user query, understanding module recognize the potential task name from it with the predefined supported tasks as additional input. If users ask for the out-of-scope tasks which has not been supported in the UniEnv, we will directly recommend user to select task from the supported list.
@@ 2.2.2. Task Planning Module
Follow the task understanding module is the planning module which is designed to generate the subsequent execution plan for efficient experiments of trajectory modelling. The input of the planning module is the task name and description from the undertanding module, the supported data and model with brief description from UniEnv. With the carefully designed prompt, the generated execution plan will contain the data name and model name for the given task, and also the detailed model optimization plan.
Due to the characteristics of different tasks and existing practice, not all the model optimization are necessary for each task. If possible, skipping the optimization step which is time-consuming and costly can accelerate the whole procedure without sacrificing performance. For example, while the data augmentation is very important for sparse check-in trajectory data, it is not so useful for dense GPS trajectory in many applications. Thus, we can generate simplified execution plan for dense GPS trajectory in some tasks to improve the success rate of the whole task and also reduce the cost. After generating the plan, it will start a simple execution step to verify the feasibility of the plan. Once any error occurs during the execution, e.g., the model name is wrong, the planning module will obtain the feedback from UniEnv and start to regenerate a new plan with the last plan as the failed history in its memory.
@@ 2.2.3. Task Execution Module
Give the execution plan, the task execution module is responsible for invoking UniEnv to execute the experiment plan. In addition to the previously mentioned basic execution interface, another interface of this module is to call AutOpt module to complete the model optimization according to the optimization plan. For both interface, task execution module will give the feedback including error information for failed cases and performance metrics for success cases. Detailed designs of AutOpt are introduced in the following section 2.3.
@@ 2.2.4. Task Summary Module
2.3. AutOpt

### v2

# TrajAgent: An LLM-based Agent Framework for Automated Trajectory Modeling via Collaboration of Large and Small Models

@@ Abstract.
Trajectory modeling, which includes research on trajectory data pattern mining and future prediction, has widespread applications in areas such as life services, urban transportation, and public administration. Numerous methods have been proposed to address specific problems within trajectory modeling. However, the heterogeneity of data and the diversity of trajectory tasks make effective and reliable trajectory modeling an important yet highly challenging endeavor, even for domain experts. In this paper, we propose TrajAgent, a agent framework powered by large language models (LLMs), designed to facilitate robust and efficient trajectory modeling through automation modeling. This framework leverages and optimizes diverse specialized models to address various trajectory modeling tasks across different datasets effectively.  In TrajAgent, we first develop UniEnv, an execution environment with a unified data and model interface, to support the execution and training of various models. Building on UniEnv, we introduce an agentic workflow designed for automatic trajectory modeling across various trajectory tasks and data. Furthermore, we introduce collaborative learning schema between LLM-based agents and small speciallized models, to enhance the performance of the whole framework effectively. Extensive experiments on four tasks using four real-world datasets demonstrate the effectiveness of TrajAgent in automated trajectory modeling, achieving a performance improvement of 2.38%-34.96% over baseline methods.
@@
copyright: acmcopyrightcopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Figure 1. The paradigm of LLM based automated trajectory modeling framework TrajAgent.
@@
With the rapid development of web services and mobile devices (Zheng, 2015; Chen et al., 2024), large-scale trajectory data, such as check-in data from social network (Yang et al., 2016), have been collected, greatly facilitating research in trajectory modeling. Trajectory modeling involves the processing, mining and prediction of trajectory data, with widespread applications in urban transportation, location services and public management. The typical areas of trajectory modeling (Chen et al., 2024; Graser et al., 2024) can be classified into five main categories: trajectory representation (Gong et al., 2023), trajectory classification (Liang et al., 2022), trajectory prediction (Yang et al., 2022), trajectory recovery (Si et al., 2023), and trajectory generation (Wei et al., 2024). Each category encompasses various sub-tasks; for instance, the trajectory prediction task can be further divided into next location prediction task (Liu et al., 2016), final destination prediction task (Zhang et al., 2019), and travel time estimation task (Wang et al., 2018), among others. Given the huge value of trajectory modeling in diverse practical applications, various algorithms and models (Graser et al., 2024) have been proposed to address these tasks, particularly deep learning-based models in recent years. This has facilitated significant advancements in the field, with many tasks achieving a high level of performance.
@@
However, existing methods are designed for specific tasks and datasets, making it difficult to share them across different tasks and data sources. For example, TrajFormer (Liang et al., 2022) is tailored for trajectory classification and cannot be applied in trajectory prediction or trajectory generation. Flash-back (Yang et al., 2020) is designed for sparse check-in trajectory prediction and is not suitable for dense GPS trajectory or road network-based trajectory modeling. In other words, due to the heterogeneity of application scenario and the diverse nature of trajectory data–varying in resolution, format and geographical regions–existing methods can only be applied in limited task with specific data for specific regions. While some early studies have explored effective trajectory modeling via unified framework (Lin et al., 2024; Zhu et al., 2024; Wu et al., 2024a), they often face several limitations: 1) their performance on individual tasks lags behind that of specialized models; 2) the range of supported trajectory modeling tasks remains limited; and 3) their training and inference processes are complex and non-trivial. Thus, despite their significant contributions, there is still a long way to go, necessitating further exploration of more effective and reliable trajectory modeling frameworks.
In recent years, the rapid development of large language models (LLMs) (OpenAI, 2024; Touvron et al., 2023) with extensive commonsense and powerful reasoning abilities presents enable the LLM based agent (Xi et al., 2023; Sumers et al., 2023) as a new paradigm for solving complex task, such as automated software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024) and automated machine learning tasks (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024b). For example, HuggingGPT (Shen et al., 2024) utilizes LLM as a core manager to address various machine learning tasks with existing AI models, VisionLLM (Wu et al., 2024b) investigates unified modeling for vision tasks across different vision domains. Inspired by these, we explore the potential of leveraging an LLM-based agent framework for automated trajectory modeling, paving the way toward effective and reliable trajectory modeling solution. Specifically, our approach seeks to harness the capabilities of LLMs to establish a collaborative framework between LLMs and various specialized models, enabling the automated and unified trajectory modeling. However, designing such an LLM-based agent for this purpose presents several significant challenges. Firstly, how to handle and integrate the diverse trajectory data and specialized models for different trajectory modeling tasks into a single, unified framework is non-trivial. Secondly, the numerous steps involved in transforming raw trajectory data into the final model output are lengthy and cumbersome (Chen et al., 2024), making full automation of the process difficult and leading to a large action space for both planning and execution. Finally, while model performance heavily depends on delicate and specialized data adaptation and model optimization, the ultimate challenge lies in effectively automating the optimization of these adaptation processes.
@@
In this paper, we propose TrajAgent, a systematic agent framework for automated trajectory modeling across diverse tasks and data. First, we design a unified environment, UniEnv, to process diverse trajectory data and provide a cohesive runtime environment for various trajectory modeling tasks within TrajAgent. In UniEnv, we define unified data and model interfaces to facilitate the seamless execution of different trajectory modeling methods. Building upon this environment, we develop an agentic workflow within TrajAgent for the automatic multi-step planning and execution of trajectory modeling tasks. The diverse trajectory modelling task workflow is decomposed into four unified steps: task understanding, task planning, task execution, and task summarization. For each step, we design an expert agent to perform the corresponding operations effectively. Finally, we introduce a collaborative learning schema that integrates agent learning through reasoning with model learning through training, enabling the effective optimization of model performance for specific data and tasks. In summary, our contributions are as follows,
@@
  • To the best of our knowledge, TrajAgent is the first LLM-based agent framework for automated and unified trajectory modeling across diverse data and tasks. It decomposes the trajectory modeling process into several sub-tasks, with expert agents designed to each.
  • To support the automated execution of various trajectory modeling tasks, we provide an unified running environment UniEnv by integrating diverse trajectory data and specialized models.
  • Furthermore, we develop a collaborative learning schema between high-level agents and low-level models to effectively enhance the final performance of proposed framework on targeted data and tasks.
@@
  • Extensive experiments on four representative trajectory modeling tasks across four real-world datasets demonstrate the effectiveness of the proposed framework, with the optimized model achieving a performance gain of 2.38% to 34.96% over baseline methods.
@@ 2. Methods
Figure 2. The framework of TrajAgent is depicted in the center, consisting of four stages: task understanding, task planning, task execution, and task summary. Besides, to ensure smooth operation across diverse data and tasks, we introduce UniEnv, an experimental platform with a unified data and model interface, shown on the left side of the figure. Finally, on the right side of the figure, we present the collaborative learning schema in TrajAgent, which includes high-level learning from knowledge reasoning for agents and low-level learning from data training for specialized models.
2.1. Overview of TrajAgent
Figure 1 presents the comparison between our work and existing works. Over the past few decades, researchers have developed various task-specific models for solving single tasks on limited datasets, as shown in the top-left of Figure 1. Recently, some works (Zhu et al., 2024; Lin et al., 2024; Wu et al., 2024a) have explored the potential of building unified models for multiple trajectory modeling tasks, as shown in the top-right of Figure 1. However, limitations in available data and the diversity of tasks have constrained these unified models to a narrow range of tasks. Additionally, these models are not easy to train and utilize, making optimization challenging and far from being simple and user-friendly. Thus, following the success of LLM-based agentic framework in other domains, we propose to build a unify framework to enable the automated trajectory modeling via the collaboration between LLM-based agents and smaller specialized models. The framework is presented in the bottom of Figure 1. In this framework, the LLM serves as a controller and processor, coordinating with specific smaller models to accomplish specific trajectory tasks. This approach enables users to effortlessly extend support for various data and models without delving into the intricate details of individual models. In essence, users can view the entire framework as a unified modeling platform, achieved through automated trajectory modeling across a variety of tasks and data. As TrajAgent A is to generate an optimized models M’, based on a user query q, optional trajectory data D, and the selected raw model M, the task definition is as follows,
$M^{\prime}=\arg\min_{M}\mathcal{L}(M,T,q,D,A),$
$\text{where }T=A(q),D=A(q,T),M=A(T,D).$
Figure 2 presents the whole framework of TrajAgent. It contains three key components: (1) UniEnv, an environment with a unified data and model interface that supports the execution and training of various trajectory models; 2) Agentic Workflow, which is designed to automatically decompose and complete diverse trajectory modelling tasks; 3) Collaborative Learning Schema, an additional automated optimization module for enhancing the performance of specific model through the collaboration between LLM-based agents and specialized models with different learning mechanisms. Details of each component are presented as follows.
2.2. UniEnv: Environment for Experiments
@@ 2.2.1. Task Interface
Figure 3 summarizes the trajectory modeling tasks and associated models supported by UniEnv. The framework covers 5 fundamental trajectory modeling tasks: prediction, recovery, classification, generation, and representation, utilizing a total of 18 methods. For tasks such as prediction, recovery, and classification, we further introduce several subtasks. For example, the prediction task includes subtasks like next location prediction and travel time estimation, the classification task includes subtasks like trajectory user linking, intention prediction and anomaly detection. To enhance the clarity and effectiveness of understanding users’ language queries, we provide a detailed language description for each task. This description helps extract precise task requirements from user queries and facilitates subsequent data and model selection processes.
@@ 2.2.3. Model Interface
As previously mentioned, we support 18 models across 5 fundamental trajectory modeling tasks. To support training these models in TrajAgent, we select at least one well-known model for each task and adapt them to match the running environments in UniEnv. Furthermore, we extract the semantic context from original paper of each model with the txyz.ai APIs  [^1] https://www.txyz.ai/ to generate detailed description of model. In this description, the verified data information and supported trajectory task information are recorded which supports the data and model selection in the subsequent trajectory planning stage of TrajAgent.
@@ 2.2.4. External Tools
To extend the capabilities of TrajAgent, we collect several external tools in UniEnv, including paper context extraction tool txyz.ai, hyperparameter optimization tool optuna  [^2] https://optuna.org/, processing and visualization tool for trajectory data movingpandas  [^3] https://github.com/movingpandas/movingpandas, open street map data processing tool osmnx  [^4] https://github.com/gboeing/osmnx. Here, we also regard the LLM APIs utilized in the agentic workflow as one of the interface in the UniEnv, including the ChatGPT API  [^5] https://openai.com/ and DeepInfra  [^6] https://deepinfra.com for open-source LLMs.
Figure 3. The TrajAgent framework supports 5 fundamental trajectory modeling tasks, encompassing a total of 18 methods. These tasks include trajectory prediction, recovery, classification, generation, and representation. For the first three tasks, several common subtasks are also provided. Each task is equipped with at least one method, ensuring comprehensive coverage of foundational trajectory modeling objectives and SOTA algorithms. Detailed introduction of methods can refer to the appendix A.
2.3. Agentic Workflow of TrajAgent
As shown in the middle of Figure 2, the agentic workflow of TrajAgent is organized into four key modules: task understanding, task planning, task optimization, and task summary, which form an automated processing chain from user query to final result, eliminating the need for human-in-the-loop. Specifically, the task understanding module first receives user instructions in natural language form, and analyzes and identifies the type, name, and other key information of the tasks involved. Then, the task planning module will plan for the identified task, including dataset matching and model selection. Next, the task execution module executes the planning task and cooperate with the additional performance optimization module collaborative learning schema to further improve the task performance from both the agent learning and model learning perspectives. Last, the task summary module generate an analytical report of the task based on historical interactions and decisions of TrajAgent. Following the common practice(Shinn et al., 2024), each module in TrajAgent can be regarded as a small agent, which consists of a function module for executing its core function, a memory for recording the history interaction, and a reflection module for learning practice from the memory.
@@ 2.3.1. Task Understanding Module
As the first module of TrajAgent workflow, task understanding module is designed to interact with user and extract detailed task information to launch subsequent stages. Given the user query, understanding module recognize the potential task name from it with the predefined supported tasks as additional input. If users ask for the out-of-scope tasks which has not been supported in the UniEnv, we will directly recommend user to select task from the supported list.
@@ 2.3.2. Task Planning Module
Follow the task understanding module is the planning module which is designed to generate the subsequent execution plan for efficient experiments of trajectory modelling. The input of the planning module is the task name and description from the understanding module, the supported data and model with brief description from UniEnv. With the carefully designed prompt, the generated execution plan will contain the data name and model name for the given task, and also the detailed model optimization plan. Due to the characteristics of different tasks and existing practice, not all the model optimization are necessary for each task. If possible, skipping the optimization step which is time-consuming and costly can accelerate the whole procedure without sacrificing performance. For example, while the data augmentation is very important for sparse check-in trajectory data, it is not so useful for dense GPS trajectory in many applications. Thus, we can generate simplified execution plan for dense GPS trajectory in some tasks to improve the success rate of the whole task and also reduce the cost. After generating the plan, it will start a simple execution step to verify the feasibility of the plan. Once any error occurs during the execution, e.g., the model name is wrong, the planning module will obtain the feedback from UniEnv and start to regenerate a new plan with the last plan as the failed history in its memory.
@@ 2.3.3. Task Execution Module
Give the execution plan, the task execution module is responsible for invoking UniEnv to execute the experiment plan. In addition to the previously mentioned basic execution interface, another interface of this module is to call collaborative learning schema module to complete the model optimization automatically. For both interface, the task execution module will give the feedback including error information for failed cases and performance metrics for success cases.
@@ 2.3.4. Task Summary Module
2.4. Collaborative Learning Schema
Due to the geospatial heterogeneity and the diversity of trajectory data, the trajectory models usually cannot be directly transferred between data and regions. In other words, for various data in different region, the model needs to be trained from scratch. Thus, the sufficient optimization of various models with targeted data becomes emergent. In TrajAgent, we design collaborative learning schema to complete this automatic optimization and generate optimized specialized models for targeted task and data. As shown in the right part of Figure 2, the collaborative learning framework involves two levels of learning: high-level knowledge reasoning for agents and low-level data training for specialized models. The high-level agent proposes training settings for the low-level models based on its expert knowledge and experimental records. The low-level models are then trained using these settings, and their performance metrics are reported back to the high-level agent for further collaborative learning. This iterative process continues until the performance meets the predefined requirements or the maximum number of exploration epochs is reached.
2.4.1. Agent Learning via Reasoning